OpenStudy (anonymous):

I need to rewrite the constraints in slope in interception form for 4x=3y>30,x+3>30,x>0 y=70

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (ranga):

4x=3y>30 does not look right. Is = supposed to be + ?

OpenStudy (anonymous):

NO, I THINK IT'S GREATER THAN

OpenStudy (ranga):

You have to type the question correctly. It is not making sense.

OpenStudy (anonymous):

THERE IS AN UNDERSCORE UNDER THE >'S

OpenStudy (ranga):

Is it \[\Large 4x = 3y \ge 30\]?

OpenStudy (anonymous):

NO, ITS 4X + 3> WITH AN UNDERSCORE UNDER THE GREATER SIGN, ALL THE >'S HAVE UNDERSIGNS OM THEM, JUST DON'T KNOW HOW TO CORRECT THAT PART.

OpenStudy (ranga):

see now you are saying 4x + 3. In the question you have 4x = 3y. And my first question was: Is the equal sign supposed to be a + sign. You said No. Also now you have 4x + 3 instead of 4x + 3y. If you cannot type the question correctly how can you expect someone to help? Use the Equation button under the reply box to type what you mean.

OpenStudy (anonymous):

THE ORIGINAL PROBLEM IS 4X + 3Y> 30 X + 3Y>21 X>0 Y=70

OpenStudy (anonymous):

\[4X + 3Y \ge 30 X + 3Y \ge 21 X O =70\]

OpenStudy (anonymous):

YES. YOU HAVE THE CORRECT PROBLEM.

OpenStudy (ranga):

I am sorry. Each time you type the problem the equation keeps changing so much I just don't know what is it I am supposed to do. XO = 70 does not make sense. Also all the equations crunched up together changes the meaning. Also, the equation editor does not take spaces to separate words and so you will have to type just one line of equation at a time.

OpenStudy (anonymous):

GO UP 4 REPLIES. THAT IS THE CORRECT PROBLEM

OpenStudy (ranga):

What is XO = 70 at the end mean?

OpenStudy (ranga):

I will leave this for someone else to answer. Sorry I can't be of help.

OpenStudy (anonymous):

SORRY. THAT'S NOT A 70. ITS A > SIGN WITH A 0.